68$ and 25cent divided by what equals 4 dollar and 50 cent?
Answer by checkley77(12844)

68.25/x=4.50 cross multiply
4.50x=68.25
x=68.25/4.50
x=91/6
Proof:
68.25/(91/6)=4.50
4.5*91/6=68.25
68,25=68.25